OpenStudy (anonymous):

What is the equation of the line in slope-intercept form that passes through the point (3, –1) and is perpendicular to the line y = x + 1?

OpenStudy (anonymous):

If it's perpendicular then it has a negative reciprocal for the slope. m = -1 -1 = -1(3) + b -1 = -3 + b 2 = b y = -x + 2